PNG electric company manufactures a number of electric products. Rechargeable light is one of the PNG’s products that sells for $180/unit. Total fixed expenses related to rechargeable electric light are $270,000 per month and variable expenses involved in manufacturing this product are $126 per unit. Monthly sales are 8,000 rechargeable lights.

If PNG  decreases selling price by 10% and sells 10,000 units what is the profit or loss?

200,000

90,000

-50,000

300,00
Answer #1

Calculation of profit/loss:

Selling price after decrease= 180*(1-0.10)= $162

Variable cost per unit= $126

Contribution= selling price-variable cost=162-126= $36

Contribution for 10000 units= 36*10000= 360000

Profit= contribution-fixed cost= 360000-270000= 90000

So correct answer is 90000